Work out the specific heat capacity of a substance if it takes 5200J of energy to heat 2kg of it from 42°C to 47°C.
1
Expert's answer
2020-06-18T10:58:33-0400

We can find the specific heat capacity of a substance from the formula:


"Q = mc \Delta t,""c = \dfrac{Q}{m \Delta t},""c = \dfrac{5200 \ J}{2 \ kg \cdot (47 ^{\circ}C - 42 ^{\circ}C)} = 520 \ \dfrac{J}{kg \cdot \! ^{\circ}C}."

Answer:

"c = 520 \ \dfrac{J}{kg \cdot \! ^{\circ}C}."